OpenStudy (anonymous):

A construction company rented bins from a waste collection agency for a large-scale clean-up. The bins are in the shape of rectangular prisms. The dimensions of a bin are shown below. • Length – 8 feet • Width – 6 feet • Height – 5.5 feet How much waste material did the construction company collect if 7 such bins were filled to the brim with waste material? 1,848 cubic feet 1,078 cubic feet 10,164 cubic feet 1,750 cubic feet

OpenStudy (anonymous):

This is a volume question. V = LWH = 8*6*5.5 = 264. There are 7 such bins so 264*7 = 1848. 1848 is your answer.

OpenStudy (anonymous):

Volume of 1 bin = l*b*h = 8*6*5.5 = 264 cu ft so 7 full containers = 7*264 cu ft = 1848 cu ft